What is the concentration of Br- in a solution prepared by mixing 0.500 L of 0.0450 M KBr and 0.250 L of 0.0250 M CaBr2?


1
Expert's answer
2022-05-24T06:47:03-0400

Solution:

Moles of solute = Molarity × Liters of solution

Therefore,

Moles of KBr = (0.0450 M) × (0.500 L) = 0.0225 mol

Moles of CaBr2 = (0.0250 M) × (0.250 L) = 0.00625 mol


KBr and CaBr2 are strong electrolytes that completely dissociate in solution.

KBr(aq) → K+(aq) + Br(aq)

According to stoichiometry: Moles of Br = Moles of KBr = 0.0225 mol

CaBr2(aq) → Ca2+(aq) + 2Br(aq)

According to stoichiometry: Moles of Br = 2 × Moles of CaBr2 = 2 × 0.00625 mol = 0.0125 mol


Total number of Brmoles = (0.0225 mol + 0.0125 mol) = 0.0350 mol

Total liters of solution = (0.500 L + 0.250 L) = 0.750 L

Thus,

Concentration of Br = (0.035 mol / 0.750 L) = 0.0467 M


Answer: The concentration of Brin a solution is 0.0467 M